OpenStudy (anonymous):

Dan wants to save $11,000 for college tuition in 6 years. How much money should he put into the account that earns 6.5% interest compounded continuously?

OpenStudy (anonymous):

185

OpenStudy (anonymous):

\[P(t)=P_{0}e^{kt}\] \[P_{0}=\frac{P(t)}{e^{kt}}=\frac{11000}{e^{.065*6}}=$7447.63\]